x+ 8y = 5
x = 4-8y

Answer by richwmiller(17219)

plug 4-8y in for x in the first equation and solve for y
4-8y+8y=5
4=5
There is no solution.
They are parallel.
rearrange the second and we get
x+8y=4
x+8y=5